OpenStudy (anonymous):

Help with this Please Each statement describes a transformation of the graph of y = x. Which statement correctly describes the graph of y = x + 7? A. It is the graph of y = x where the slope is increased by 7. B. It is the graph of y = x translated 7 units to the right. C. It is the graph of y = x translated 7 units down. D. It is the graph of y = x translated 7 units to the left.

ganeshie8 (ganeshie8):

k units left transformation : f(x) -> f(x + k)

OpenStudy (anonymous):

so it would be the opposite of what i was thinking it would actually be D

ganeshie8 (ganeshie8):

Yes ! D is right, more on this here : http://www.mathsisfun.com/sets/function-transformations.html

OpenStudy (anonymous):

Express the terms of the following sequence by giving an explicit formula. 5 , 3 , 1 , -1, -3 , . . . A. \[^an\] = n + 2, where n = 1, 2, 3, 4, . . . B. \[^an\] = 2n, where n = 1, 2, 3, 4, . . . C. \[^an\] = 7 - 2n, where n = 1, 2, 3, 4, . . . D. \[^an\] = 2 + 7n, where n = 1, 2, 3, 4, . . .

ganeshie8 (ganeshie8):

see any pattern in the given sequence : 5 , 3 , 1 , -1, -3 , . . . ?

OpenStudy (anonymous):

yes it actually looks like a mirror

ganeshie8 (ganeshie8):

and.. how does each next term is increasing/decreasing by ?

ganeshie8 (ganeshie8):

looks each term is getting reduced by 2, eh ? 5-2 = 3 3-2 = 1 1-2 = -1 -1-2 = -3 ...

OpenStudy (anonymous):

hmmm so wouldnt tht make this C

ganeshie8 (ganeshie8):

Yep ! first term = 5 common difference = -2 so the nth term is ; \(a_n = a_1 + (n-1)*d\) \(a_n = 5 + (n-1)*(-2)\)

ganeshie8 (ganeshie8):

simplifying it gives u C^